OpenStudy (minthia):

What is the slope of the line that passes through the given points 2,12 and 6,11

OpenStudy (theopenstudyowl):

m = y2-y1/x2-x1 = 11-12/6-2 = -1/4; slope = m = -1/4

OpenStudy (minthia):

thanks